OpenStudy (anonymous):

A man walks on a straight road from his home to market 2.5kn away with a speed of 5km/hr . finding the market closed he instantly walks back home with a speed of 7.5km/hr . find the magnitude for avg speed and avg velocity over the following time (a) 0 to 30 minutes. (b) 0 to 50 minutes. (c) 0 to 40 minutes. ( With each and every step if possible. )

OpenStudy (anonymous):

@anusha.p can you help me with this sum .. please ?

OpenStudy (anonymous):

i'm very bad at physics.. i'm sorry...@ashna

OpenStudy (anonymous):

its okay .. btw tnx :)

OpenStudy (anonymous):

avg speed= total distance covered/ time taken avg velocity= total displacement/time taken  a) In the first half hour, the man covers 2.5 km(his speed is 5km/h). So the avg speed is 5km/h(given). Or use the above formula. Distance = 2.5km. Time is 1/2h The avg velocity is also the same since distance and displacement are equal. b) For the first 50 min -For the first 30 min distance covered is 2.5 km -For the next 20 min(1/3 hours), the man covers a distance of 7.5*1/3=2.5 km. So the total distance covered= 2.5+2.5 time taken=50 min=5/6 hrs Avg speed=6km/h Here the avg velocity will be different since distance and displacement are not equal. Here,the distance covered in the last 20min at 7.5km/h will get subtracted from the distance covered in the first 30min to get displacement(which is 0).So the avg velocity is 0.